OpenStudy (anonymous):

Hi can someone please tell me how i can show that f=mv^2/R may be written as F=mw^2R

OpenStudy (jamesj):

The arc length of a circle of radius R and angle \( theta \) is \[ \theta R \] Hence if a mass moves an angular distance \( \Delta \theta \) in time \( \Delta t \), it has velocity \[ v = \frac{\Delta \theta}{\Delta t} R = \omega R \] by definition of angular velocity \( \omega \). Now substitute this expression into your first one to get the second.

OpenStudy (anonymous):

im just a little confused bc in my book it says w= theta/time

OpenStudy (jamesj):

Yes, \[ \omega = \frac{d\omega}{dt} \] the rate of change of angle. That is, \( \omega \) is the angular velocity. For example, if the angular velocity \( \omega = 2\pi \), then every second, the circle/disk makes one complete revolution of \( 2\pi \) radians.

OpenStudy (jamesj):

In that case, if you the circle had radius r, a point on the circle traveled a distance of \[ 2\pi r \] in that one second and hence it's average velocity has magnitude \[ v = \omega r = 2 \pi r \] also.

OpenStudy (jamesj):

*average instantaneous velocity v. It's average velocity of course is zero. But at any given moment its instantaneous velocity or speed is \( \omega r \).

OpenStudy (anonymous):

can u please show how you got v=dellta theta/change time R=wR step by step bc i am a little stuck

OpenStudy (jamesj):

Let's come at this a different way. Suppose an object is in uniform circular motion of radius R and angular velocity \( \omega \). Over a very small time period \( \Delta t \), what is the arc length the object has moved?

OpenStudy (jamesj):

To answer that question, first tell what angle the object has moved.

OpenStudy (anonymous):

arc length= R(radius)w ( angular velocity)

OpenStudy (jamesj):

No. What is the angle the object moves in time \( \Delta t \).

OpenStudy (jamesj):

If the object is moving with angular speed w (for omega), then in time Dt (for delta t), the **angle** it has swept out is theta = w.Dt make sense?

OpenStudy (jamesj):

because that is what angular velocity omega means. By definition, w is the amount of angle moved in 1 second. Hence in Dt seconds, the object must have moved an angle of \[ \theta = \omega \Delta t \] Yes?

OpenStudy (anonymous):

yah

OpenStudy (jamesj):

Hence in time Dt, what **distance** has the object moved?

OpenStudy (anonymous):

arc length= Rxw not sure

OpenStudy (jamesj):

If it has moved an angle of theta, how far in **distance** has it moved?

OpenStudy (jamesj):

what is the arc length for the angle theta, in other words?

OpenStudy (anonymous):

other formulas that were given with this problem were v=2piRn/deltat , w=2pin/delta t and w= Dtheta/ delta t.. this problem is like a warm up to the chapter i was wondering can i use this formulas to derive it or do i have to use your method

OpenStudy (jamesj):

In time \( \Delta t \) the angle moved is \[ \Delta \theta = \omega \Delta t \] Hence in that time, \( \Delta t \), the distance moved is \[ \Delta s = \Delta \theta . R \] I'm taking you back to just basic, basic definition here.

OpenStudy (jamesj):

Now what is \( \Delta \theta \), we know that \[ \Delta \theta = \omega \Delta t \] Hence \[ \Delta s = \Delta \theta R = \omega \Delta t . R \]

OpenStudy (jamesj):

Divide both sides by delta t and we have \[ \frac{\Delta s }{\Delta t} = \omega R \]

OpenStudy (jamesj):

So far so good?

OpenStudy (anonymous):

yah much better

OpenStudy (jamesj):

Now what is delta s / delta t equal to?

OpenStudy (anonymous):

isnt wR

OpenStudy (jamesj):

speed v is \[ v = \frac{\Delta s}{\Delta t} \] therefore \[ v = \omega R \]

OpenStudy (anonymous):

can i use any of the formulas listed above as well to derive thisquation

OpenStudy (anonymous):

v=2piRn/delta t , w=2pin/delta t and w= Delta theta/ delta t..

OpenStudy (jamesj):

These have unusual notation. What is n here, frequency?

OpenStudy (anonymous):

R radius , n revolutions

OpenStudy (underhill):

(James, how do you do fractions in the equation editor? I'd like to volunteer an explanation.)

OpenStudy (jamesj):

Yes, frequency. In short, yes, you can use these equations to get to the same results. But in an important way, the method I have shown you is more fundamental. As an exercise for yourself, use your equations to find the same result.

OpenStudy (jamesj):

@Underhill, go for it. Fraction are written as \frac{a}{b} \[ \frac{a}{b} \]

OpenStudy (jamesj):

*fractions

OpenStudy (underhill):

Awesome!

OpenStudy (anonymous):

can someone please show how i can derive it using the formulas.. having a very difficult time

OpenStudy (underhill):

\[F = m\frac{v^2}{R}\]Since\[v = \omega \times R = \frac{radians}{second}\times\frac{meters}{radian}\]then \[v^2 = \omega^2R^2\]and therefore\[F = m\frac{v^2}{R} = m\frac{\omega^2R^2}{R} = m \omega^2R\]

OpenStudy (underhill):

It's no more complicated than this. Omega is the angular velocity. V is the linear velocity. You can already see how closely related these two equations are. You can make the connection by realizing that omega is radians/second, and that in any circle that the radius corresponds exactly to one radian. Think of the radius not as a simple distance, but in terms of "meters per radian" because that is exactly what it is. Multiplying omega (radians/second) by the radius (meters/radian) gives meters/second. Viola! Meters/second = velocity. You're there!

OpenStudy (underhill):